OpenStudy (anonymous):

How to determine if triangle ABC and triangle DEF are congruent using coordinates. (Geometry)

OpenStudy (anonymous):

I already drew the two triangles on the graph.

OpenStudy (maheshmeghwal9):

So u mean, u r given only co-ordinates. Then what should u do is this: - Apply Distance formula; i.e., \[\text{Distance}=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}.\] And prove all the sides equal by SSS congruency. Thats what u should do:)

OpenStudy (maheshmeghwal9):

Any doubt?

OpenStudy (anonymous):

let me try.

OpenStudy (maheshmeghwal9):

k:)

OpenStudy (anonymous):

okay, for AB I got \[\sqrt{36}\]

OpenStudy (maheshmeghwal9):

so calculate other sides too.

OpenStudy (anonymous):

alright

OpenStudy (maheshmeghwal9):

:)

OpenStudy (anonymous):

for the other two, I got \[\sqrt{64}\] and \[\sqrt{100}\]

OpenStudy (maheshmeghwal9):

k! just do the same for other triangle. & compare corresponding sides.

OpenStudy (maheshmeghwal9):

u must find that corresponding sides of 2 triangles are equal

OpenStudy (maheshmeghwal9):

clear?

OpenStudy (maheshmeghwal9):

????????

OpenStudy (anonymous):

okay, alright.

OpenStudy (maheshmeghwal9):

hmm........ well done:)

OpenStudy (anonymous):

all but two sides were congruent.

OpenStudy (anonymous):

all but one*

OpenStudy (anonymous):

????

OpenStudy (maheshmeghwal9):

u mean one-one side of each triangle isn't congruent?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

AC and DF

OpenStudy (maheshmeghwal9):

so must must must be a calculation mistake

OpenStudy (anonymous):

Okay, I got it. Sorry. Now they all are congruent.

OpenStudy (maheshmeghwal9):

yeah well done :D

OpenStudy (anonymous):

thanks.

OpenStudy (maheshmeghwal9):

^_^